OpenStudy (anonymous):

If a = -7, what is 3-2(4+a)³ ?

Nnesha (nnesha):

substitute a for -7

OpenStudy (anonymous):

Yeah I did that

Nnesha (nnesha):

so what did you get ?

OpenStudy (anonymous):

Do I have to add what's inside the parenthesis when I substitute the -7 in ? @Nnesha

Nnesha (nnesha):

yes right PEMDAs rule first solve whats in the parentheses and then take cube

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

Okay so I did that and I got -3

Nnesha (nnesha):

is it 3-2 front of the parentheses ?

Nnesha (nnesha):

4-7=-3 (-3)^3 <--solve that

OpenStudy (anonymous):

Wait no the 3-2 is not in the parenthesis its on the outside to the left

OpenStudy (anonymous):

Okay so I solved -3³ and I got +27

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

no don't forget the parentheses it's (-3)^3 not -3^3

OpenStudy (anonymous):

Yeah the exponent

OpenStudy (anonymous):

I already solved it

Nnesha (nnesha):

what did you get ?

OpenStudy (anonymous):

I got 27 when I solved (-3³) then I subtracted 3-2 which was on the left so now I got 28 I'm not sure if its correct

Nnesha (nnesha):

no

Nnesha (nnesha):

\[\huge\rm 3-2(4+a)^3\] substitute a for -7 \[\huge\rm 3-2(4+(-7))^3\] \[\huge\rm 3-2(-3)^3\] it is (-3) to the 3rd power not -(3^3) totally different

Nnesha (nnesha):

(-3)^3= -3 times -3 times -3 -3^3=-3 times 3 times 3 sign difference

OpenStudy (madhu.mukherjee.946):

i guess the answer is 57

OpenStudy (madhu.mukherjee.946):

coz 3-2(-3)^3 3-2(-27) 3+54 57

OpenStudy (madhu.mukherjee.946):

@Nnesha am i correct?

Nnesha (nnesha):

You already know.

OpenStudy (anonymous):

Wait so I dont have to subract 3-2 ?

OpenStudy (anonymous):

Wait never mind I got it !! :) thanks @Nnesha & @madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

no

Nnesha (nnesha):

np

OpenStudy (anonymous):

Can you help me with one more problem ?

OpenStudy (madhu.mukherjee.946):

say

OpenStudy (anonymous):

@madhu.mukherjee.946 I'm done with math for today my brain hurts 😞 tomorrow I'll finish them

OpenStudy (madhu.mukherjee.946):

okay:)